OpenStudy (anonymous):

need to solve rational expression 1+1/x-1=30/x^2+1 Can you help solve this equation

OpenStudy (anonymous):

cross multiply. it is a cubic equation

OpenStudy (anonymous):

x/x-1= 30/x^2+1 x^3+x= 30x - 30 x^3 -29x +30=0

OpenStudy (radar):

Is the equation stated correctly? Is it really: \[1+1/(x-1)=30/(x ^{2}+1)\] If so it appears the first 1 was ignored in the solution unless 1+1/(x-1) is the same as x/(x-1). It isn't

OpenStudy (radar):

I can see I was wrong about that! Never mind it is the same.